# SignCLIP: Connecting Text and Sign Language by Contrastive Learning

### v1

@@ 7 Conclusion: Where Are We for SLP?
SignCLIP demonstrates excellent in-domain performance but falls short of immediate zero-shot prediction on downstream ISLR tasks. This finding is consistent with previous CV studies (Li et al., 2017) before CLIP reached its scale. Similar models trained on smaller datasets close to an ImageNet scale performed much worse than supervised baselines on common benchmarks. However, a supervised approach usually requires meticulous, task-specific data collection. This is more demanding for SLP, a niche domain lacking human experts.
As a middle ground between full zero-shot prediction and full supervision, few-shot learning or fine-tuning is essential to tackle domain shift and is more realistic given the present methodology and data scale. The cross-lingual transfer effect is prospective for sign language due to iconicity and smaller grammatical discrepancies. With no surprise, under a sensible data condition, the universal pretraining paradigm that is transforming NLP and CV is also a promising research direction for SLP.
@@ 8 Limitations
First, to be not dataset/task/language-specific while possibly large-scale, we choose Spreadthesign as our pretraining dataset (§5.1), which is indeed highly multilingual and untied to any SLP task. However, we cannot release the dataset publicly, and future researchers who are interested in it have to first resolve or purchase the license from Spreadthesign. Fortunately, it is possible to augment or replace Spreadthesign with recently released public datasets of comparable or larger size, which should increase data density and variation sign-wise, but with potentially decreased diversity and balance among different sign languages.
Secondly, training large models on video data is very costly. The principle of this work is to finish every training process in less than three days on a single Nvidia A100 GPU, and we managed this goal (Table 4) by using pose-based models, a moderate context length of 256 frames, and limiting spoken language to English only. To scale up, multiple GPU training can be exploited, and architectural modifications that reduce the sequence length must be employed, for which we refer to several techniques discussed in §2.

### v2

@@ 8 Conclusion: Where Are We for SLP?
SignCLIP demonstrates excellent in-domain performance but falls short of immediate zero-shot prediction on downstream ISLR tasks. This finding is consistent with previous CV studies (Li et al., 2017) before CLIP reached its scale. Similar models trained on smaller datasets close to an ImageNet scale performed much worse than supervised baselines on common benchmarks. By comprehensively evaluating downstream ASL ISLR performance (Table 5), we also intend to shed some light on data efficiency, as a fully supervised approach usually requires meticulous, task-specific data collection. This is more demanding for SLP, a niche domain lacking human experts.
As a middle ground between full zero-shot prediction and full supervision, few-shot learning or fine-tuning is essential to tackle domain shift and is more realistic given the present methodology and data scale. The cross-lingual transfer effect is prospective for sign language thanks to iconicity. With no surprise, under a sensible data condition, the universal pretraining paradigm that is transforming NLP and CV is also a promising research direction for SLP.
@@ 9 Limitations
First, to be not dataset/task/language-specific while possibly large-scale, we choose Spreadthesign as our pretraining dataset (§5.1), which is indeed highly multilingual and untied to any SLP task. However, we cannot release the dataset publicly, and future researchers who are interested in it have to first resolve or purchase the license from Spreadthesign. Fortunately, it is possible to augment or replace Spreadthesign with recently released public datasets of comparable or larger size, which should increase data density and variation sign-wise, but with potentially decreased diversity and balance among different sign languages. Although pretrained with a highly multilingual dataset, this paper focuses downstream task evaluation on ASL, one of the highest-recourse sign languages. We leave further evaluation and exploration of other sign languages to future work.
Secondly, training large models on video data is very costly. The principle of this work is to finish every training process in less than three days on a single Nvidia A100 GPU, and we managed this goal (Table 4) by using pose-based models, a moderate context length of 256 frames, and limiting spoken language to English only. An ablation study on video length would be interesting to help understand the limitations of the current static embedding approach [^16] This is also relevant to the polysemy phenomenon between signs and words, i.e., one sign can mean more than one word and one word can also correspond to more than one sign.. To scale up, multiple GPU training can be exploited, and architectural modifications that reduce the sequence length must be employed, for which we refer to several techniques discussed in §2.
@@ Acknowledgments
We thank the (meta-)reviewers for their valuable feedback. We also thank Colin Leong, Gomèr Otterspeer, Jetske Adams, Oline Ranum, and Hamzah Luqman for their immediate interest and lively discussion in our work.